OpenStudy (anonymous):

The force F​ (in pounds) required to push a sofa across a floor varies directly as the weight w​ (in pounds) of the sofa. Complete parts​ (a) through​ (c) below. a. A person can push a 90​-pound sofa across a wood floor by pushing with a force of 25 pounds. Find an equation that describes the relationship between w and F. the answer is F=5/18w. then asks: How much force is required to push a 189​-pound sofa across a wood ​floor?

OpenStudy (astrophysics):

Next time put these questions in the physics section! But in any case, your problem says that \[\vec F ~~~~ \alpha ~~~ \vec W\] that means if you increase F by 100 the weight W will also increase by a 100. So the first part, lets go over it just in case, the reason \[F = \frac{ 5 }{ 18 } W\] is because the force used is 25 pounds on a weight of 90 pounds. Relating the two forces we get that the force applied is the same as 5/18 of the weight, so 25/90 = 5/18. Now you can apply this same formula to your next part.

OpenStudy (astrophysics):

You need to understand what the variable means, so weight (W) here is the 189 pound sofa